A long, straight wire of diameter 2.0mm carries a uniformly distributed current of 10A. At what distance from the axis of the wire will the magnitude of B vector be maximum.justify your answer


1
Expert's answer
2021-01-21T14:17:29-0500

For "0<r<R=2(mm)"


"\int Bdl=\mu_0I"


"B\cdot2\pi r=\mu_0\cdot \frac{I}{\pi R^2}\cdot\pi r^2\to B=\frac{\mu_0 I}{2\pi R^2}r"


For "r>R"


"B\cdot2\pi r=\mu_0 I\to B=\frac{\mu_0 I}{2\pi r}"


So, "B=B_{max}=\frac{\mu_0I}{2\pi R}" if "r=R=2(mm)" . Answer